Find the area of the triangle formed by the x- and y- axes and the line 12x + 10y = 30.
Answer by solve_for_x(190)

The x- and y- axes and the line 12x + 10y = 30 form a right triangle with the 90 degree angle at the origin.

To determine the base and height of the triangle, find the x- and y-intercepts of the given line.

To find the x-intercept, set y = 0 and solve for x:

12x + 10(0) = 30

12x = 30

x = 30/12

x = 2.5

To find the y-intercept, set x = 0 and solve for y:

12(0) + 10y = 30

10y = 30

y = 3

The base of triangle is 2.5, and the height is 3.

The area is then:

Area = (1/2)(base)(height) = (1/2)(2.5)(3) = 3.75